OpenStudy (crashonce):

In a GP with n terms, the first term is 1, the common ratio is r and the sum is s, where r and s are not zero. If each term of the GP is replaced by its reciprocal, then the sum becomes?

OpenStudy (crashonce):

@ganeshie8

OpenStudy (crashonce):

@perl

OpenStudy (zzr0ck3r):

Would it not diverge then?

OpenStudy (zzr0ck3r):

example 1/2+1/4+1/8....would become 2+4+8....?

OpenStudy (crashonce):

im not sure about that, just basic AP, GP

OpenStudy (zzr0ck3r):

I am a little out of my realm because I only took basic calc series and never studied progressions in general, but it seems from the example I showed, if I understand you correctly, that we cant guarantee it converges.

OpenStudy (crashonce):

ok

OpenStudy (crashonce):

how should i start

ganeshie8 (ganeshie8):

\[\large 1+r+r^2+\cdots + r^{n-1} = s\]

ganeshie8 (ganeshie8):

replace the left side with partial sum formula

ganeshie8 (ganeshie8):

\[\large \dfrac{1-r^n}{1-r} = s\]

OpenStudy (crashonce):

what about the reciprocals @ganeshie8

ganeshie8 (ganeshie8):

if you replace each term by their reciprocal, the series becomes: \[\large 1+\frac{1}{r}+\frac{1}{r^2}+\cdots +\frac{1}{ r^{n-1}} \] Notice that the common ratio is 1/r now

OpenStudy (crashonce):

i still dont know how to get the relation

ganeshie8 (ganeshie8):

partial sum would be \[\large \dfrac{1 - \frac{1}{r^n}}{1-\frac{1}{r}}\] rearragne a bit and try to express it in terms of \(s\)

OpenStudy (crashonce):

is it r^n/s

ganeshie8 (ganeshie8):

nope, try again

OpenStudy (crashonce):

not sure

ganeshie8 (ganeshie8):

\[\large \dfrac{1 - \frac{1}{r^n}}{1-\frac{1}{r}} \] \[\large \dfrac{(r^n - 1)/r^n}{(r-1)/r} \] \[\large \dfrac{1-r^n}{1-r} \cdot \dfrac{r}{r^n} \] \[\large s \cdot \dfrac{1}{r^{n-1}} \]

OpenStudy (crashonce):

ohh right

OpenStudy (crashonce):

thanks

OpenStudy (crashonce):

so its s/r^n-1

ganeshie8 (ganeshie8):

Yep!